Drivers A and B are 50 mi apart, and they plan to meet at 7 P.M. at a point between them that is 30 mi from A. Both drive at 50 mph. What time should each driver start out?
Answer by ewatrrr(24785)


Hi,
Note: Driver A will be driving 30mi aan Driver B will driving 20mi
d= r*t or t= d/r
Driver A
t = 30mi/50mph = 3/5 hr = 36/60 or 36 min. Driver A should leave at 7:36 P.M.
Driver B
t = 20mi/50mph = 2/5 hr = 24/60 or 24 min. Driver B should leave at 7:24 P.M.